Session: Nov 1997 (session 4)
Wavelength: 3.6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (C. Bortolotti at Sat Nov 22 19:00:12 MET 1997) Success
Project el020b - no known problems. Bad time (fog and rain)

Onsala (A. Polatidis at Sat Nov 22 19:22:32 MET 1997) Success
No known problems. Light rain for a few hours

Effelsberg (dgraham at Mon Nov 24 11:59:03 MET 1997) Minor failures
High humidity. Noise temperature high (90K) or limiting in IF, not clear excatly why.

Noto (C.Trigilio at Tue Nov 25 17:02:15 MET 1997) Minor failures
The operator reports a PROBABLE jump in the formatter (early of 1 sec) at the end of the experiment. This problem likely occurred after 15 UT.

Calibration (Fredrik T Rantakyro at Tue Dec 9 15:07:12 MET 1997) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the NOV97 directory under the name el20b.antab Read the comments in the file to get further information.